OpenStudy (anikay):

Find the point(s) (x,y) on the hyperbola y^2 = 1+(x^2/4) that is are closest to the point (6,0). Calculus. Please explain how to work the problem.

OpenStudy (anonymous):

IDK sorry

OpenStudy (anonymous):

idk i will try to figure out though

OpenStudy (anikay):

\[y^2= 1+\frac{ x^2 }{ 4 }\] just to clarify the equation and thank you for the help

OpenStudy (anonymous):

well you need to get y alone by square rooting the problem for starters

OpenStudy (anikay):

that's one way of doing it. I personally think it'd be easier just to derive it implicitly.

OpenStudy (anonymous):

yes that is one way

OpenStudy (anikay):

so \[y'2y = \frac{ x }{ 2 }\] therefore \[y' = \frac{ x }{ 4y }\]

OpenStudy (anonymous):

you are a guiness

OpenStudy (anikay):

this is where I'm lost though. I don't know how to find out the distance. I can get the critical points here but at (6,0) this derivative is infinity which shows the point clearly doesn't exist, which is fine. but I'm just lost. I'll look it up online or somethin

OpenStudy (anonymous):

im a fricken 8 year old i don't know anything!

OpenStudy (anonymous):

sike I'm actually 17:)

OpenStudy (anikay):

are you really?

OpenStudy (anikay):

oh okay xD I'm only 18 so its cool

OpenStudy (phi):

I would define a function that represents the distance from the point (6,0) to the hyperbola, and then minimize that function. It is perhaps easier to deal with the square of the distance (minimizing the distance squared will also minimize the distance) use the distance formula \[ D^2 = (x_1 - x_0)^2 + (y_1 - y_0)^2 \] use (6,0) for \( (x_0, y_0) \) and \[ y = \sqrt{1 + \frac{x^2}{4}} \]

OpenStudy (phi):

we have the problem of minimizing the function \[ f(x) = (x - 6)^2 + \left(0 - \sqrt{1+\frac{x^2}{4} }\right)^2\\ f(x) = (x - 6)^2 + 1 + \frac{x^2}{4} \]

OpenStudy (phi):

Here is a graph of the problem